Solve.
When making a telephone call using a calling card, a call lasting 5 minutes costs $1.70. A call lasting 14 minutes costs $3.95. Let y be the cost of making a call lasting x minutes using a calling card. Write a linear equation that models the cost of making a call lasting x minutes.
Answer Choices
y = 0.25x + 0.45
y = 0.25x - 10.05
y = -0.25x + 2.95
y=4x-183/10

Answer by mananth(16946)

When making a telephone call using a calling card, a call lasting 5 minutes costs $1.70.
y = 0.25x + 0.45
= 0.25+0.45 = 1.7
A call lasting 14 minutes costs $3.95.
y = 0.25x + 0.45
= 0.25*14 +0.45=3.95